# Modelling Sampling Distributions of Test Statistics with Autograd

### v1

@@ 3.2 Directly modelling the empirical cdf
Figure 3: ON/OFF cdfs from modelling the empirical cdf with the latter as the targets at different $(\mu,\nu)$ points, with the associated $68\%$ confidence band attained from conformal inference.
@@
Figure 4: ON/OFF pdfs obtained by differentiating the models of the empirical cdf at different $(\mu,\nu)$ points, with the associated $68\%$ confidence band attained from conformal inference..
@@
Trying to model an intrinsically discrete distribution with a smooth approximation may be beyond an algorithm such as ALFFI and one might argue it is in any case a hopeless task. Nevertheless it is useful to check if an alternative, more direct, approach works better: directly modelling the empirical cdf. At each $\{\mu_{i},\nu_{i}\}$ point, we generate $K=100$ experiments: $\{N_{i}^{j}\}_{j=1}^{K}\sim\text{Poiss}(\mu_{i}+\nu_{i})$ and $\{M_{i}^{j}\}_{j=1}^{K}\sim\text{Poiss}(\nu_{i})$, for which we generate $K$ instances of the test statistic $\{\lambda_{i}^{j}\}_{j=1}^{K}$ according to Eq. 10. These data are used to calculate the empirical cdf $F_{i}(\lambda_{i})$ at every parameter point. The data were then flattened to result in a training set of size $B=10^{7}$, $\mathcal{T}=\{(\mu_{i},\nu_{i},\lambda_{i},F_{i})\}_{i=1}^{B}$.
@@
The cdf was again approximated with a neural network with 3 input features $\mathbf{x}=\{\mu,\nu,\lambda\}$ but this time with the targets set to values of the empirical cdf $F$. The architecture consisted of 6 hidden layers with 12 nodes each and a single output. The activation function at each hidden node is a SiLU. The network was trained to minimize the mean square error with the NAdam optimizer with a fixed learning rate of $3\times 10^{-5}$. The training set is composed of $10^{7}$ examples, which were used in batches of size $60$ for $10^{6}$ iterations, that is, for 6 epochs.
@@
A much better approximation of the cdf is obtained, as shown in Fig. 3 and more importantly in the associated pdf. Fig. 4 shows that the jagged fluctuations in the pdf disappear and we obtain a smooth approximation of the discrete distribution. Modelling the empirical cdf presents yet another advantage: the possibility of using the empirical cdf as the true response variable in the conformal inference algorithm (see Sec. 2.3) and thereby providing a quantification of the uncertainty in the modelling of the cdf and pdf. We use conformal inference to calculate a $68\%$ confidence interval in the cdf space for each value of the test statistic. For the associated pdf, a coarse-graining of $\lambda$ is performed by histogramming $\lambda$ and the coarse-grained $f(\lambda)$ was used as the true response variable (that is, the true pdf). The excellent agreement between the predicted pdf and the histogrammed pdf is shown in Fig. 4 along with the $68\%$ conformal confidence interval at each value of $\lambda$.
3.3 Multistage Modelling of the CDF
@@
Quantile residuals ($\Delta C$): Let $p=\{F_{1},...,F_{k}\}$ be a set of evenly spaced probabilities and let $\hat{F}=\{\hat{F}_{1},...,\hat{F}_{n}\}$ be a set of estimated cdf values at a given $(\mu,\nu)$ parameter point with $n>>k$. Define $q$ as the quantiles of $\hat{F}$ associated with the probabilities $p$. If we can model the correction function $p=C(q)$ that maps the estimated cdfs $q$ to the exact cdf $p$ with sufficient accuracy then, in principle, the correction function can be used to improve the estimated cdf. Writing $p=q+\Delta C(q)$, we define the quantile residual function as $\Delta C=q-p$ and try to model it as a function of $q$ following the multistage method.
@@
Nevertheless, we wish to see whether a neural network can produce a smooth interpolation of this noisy function by implementing the suggestions described in [27]: 1. Use a sine nonlinearity in the first layer and tanh nonlinearity in subsequent layers. 2. Use Glorot weight initialization [34] with a constant factor $\kappa=60$ multiplying the weights in the first layer and no such factor in subsequent layers. It was observed that the $\kappa$ weight factor helps to stabilize the training and the sine nonlinearity helps capture the high frequency. 3. Normalize the target function by the RMS of the outputs of the target function $\Delta C\rightarrow\Delta C/RMS(\Delta C)$, as in Eq. 6. Further experiments were performed such as adding or omitting dropout, batch normalization, weight regularization, and experimenting with different nonlinear functions such as ReLU, SiLU, tanh, sine, and groupsort [35]. None of these experiments were able to capture the high frequency behavior of the residual target function when trained on the residuals of all parameters in the training data simultaneously. However, we were able to use a simple neural network to model the residual function when modeled point-by-point in the parameter space.
@@
The frequency principle [36] states that neural networks first fit low frequencies and later fit high frequencies during the training process. However, prior studies [37, 38] have demonstrated that in practice standard neural networks struggle to fit high frequency functions. We confirm these results and observe that the neural networks we used struggle to fit the residual functions even when trained for very long training runs. LSTMs, [39] on the other, hand are known to outperform standard neural networks in fitting high frequency and sequential data, where the order of the data is significant [40]. They have been successfully applied in many areas in sequence learning and time series forecasting such as financial forecasting [41], handwriting recognition [42] and speech recognition [43]. In our study, LSTMs achieve reasonably good predictions of the residuals when trained over all $(\mu,\nu)$ values. Although our LSTM models outperform our multi-layer perceptron (MLP) models in this task the results are unsuitable for use in the MSNN approach due to inadequate precision in the resulting predictions.
@@
Cdf residuals ($\Delta r$): An alternative to modelling the quantile residuals is to model the cdf residuals as a function of $\lambda$, which is closer to what is done in the MSNN approach. At a particular $(\mu,\nu)$ point, the cdf residual is simply the difference between the predicted and the empirical cdf: $\Delta r=\hat{F}-F$. Figure 10 shows that the cdf residuals as a function of $\lambda$ are a slightly smoother function of the parameter points than is $\Delta C$. Nevertheless, the MLP and LSTM models still struggle to capture the high-frequency nature of this residual function. A coarse-graining of $\lambda$ was also performed to reduced the high-frequency of the residual function but to no avail.
@@ 4 Example 2: SIR Model
The lessons learned in the ON/OFF example are applied to modelling the cdf in a different context: epidemiology. We use the SIR model [10, 11] which treats an epidemic as a system in which transitions occur between three states or compartments: susceptible (S), infected (I), and recovered (R). This model has both a stochastic realization as well as a description in terms of ordinary differential equations
@@
where $\theta=\{\alpha,\beta\}$, $x_{n}$ is the observed number of infected individuals at the observation time $t_{n}$ and $I_{n}=I(t_{n},\theta)$ is the predicted mean infected count obtained by solving the system of ordinary differential equations, Eq. 14. This example is chosen because it exemplifies the utility of the methods discussed in this paper for inference with intractable statistical models. All of our major findings from the ON/OFF example apply to this example, but here we wish to improve the modelling of the cdf and employ more techniques for uncertainty quantification.
4.1 Directly modelling the empirical cdf
@@
One of the lessons from the ON/OFF problem is that for discrete distributions the ALFFI algorithm does not yield a sufficiently accurate smooth model of the cdf, which was confirmed in the SIR example. We therefore directly model the empirical cdf as a function of the model parameters $(\alpha,\beta)$ and the test statistic. The training dataset comprises $250$ uniformly sampled $(\alpha,\beta)$ parameter points, where at each $(\alpha_{i},\beta_{i})$ point, $K=400$ epidemics were simulated yielding a set of test statistics $\{\lambda_{i}^{j}\}_{j=1}^{K}$ and a set of empirical cdfs $\{F_{i}^{j}\}_{j=1}^{K}$. The data were then flattened to result in a training set of size $B=10^{7}$, $\mathcal{T}=\{(\alpha_{i},\beta_{i},\lambda_{i},F_{i})\}_{i=1}^{B}$.
@@
The MSE loss is known to be sensitive to outliers in the data. Therefore, we chose to use the more robust Huber loss [44], which is an MSE loss for relatively small errors and an absolute loss for larger errors. The Huber loss is given by
@@
$L_{\delta}(\boldsymbol{\omega})=\left\{\begin{array}[]{ll}\frac{1}{2}(y-f(\mathbf{x},\boldsymbol{\omega}))^{2}&\text{for }|y-f(\mathbf{x},\boldsymbol{\omega})|\leq\delta\\ \delta\left(|y-f(\mathbf{x},\boldsymbol{\omega})|-\frac{1}{2}\delta\right),&\text{otherwise.}\end{array}\right.$ (16)
where $\delta$ is a tunable parameter. The Huber loss is both differentiable everywhere and robust to outliers. We employ models trained with MSE and models trained with the Huber loss and observe that for relatively short training runs ($\approx 10^{5}$ iterations) the Huber loss with $\delta=0.7$ significantly outperforms the MSE loss when all other hyperparameters of the model are held fixed. For larger training runs ($>10^{6}$ iterations), however, using the Huber loss and MSE loss yields approximately the same results. In order to reduce the number training iterations, we use the model with Huber loss for all of our subsequent models.
All of the models for this example approximate the cdf using a fully-connected neural network with 3 input features $\mathbf{x}=\{\alpha,\beta,\lambda\}$, 5 hidden layers with 10 nodes each, and a single output using a sigmoid. The activation function at each hidden node is a SiLU. The network was trained with the NAdam optimizer with a fixed learning rate of $3\times 10^{-4}$. The training set, which comprises $10^{7}$ examples, was used in batches of size $60$ for $3\times 10^{6}$ iterations, that is, for 18 epochs. As the training proceeds the network with the lowest validation loss is saved. Figures 5 and 6 show the resulting estimated cdfs and pdfs with the 68% conformal confidence intervals at different values of $\lambda$ and for different $(\alpha,\beta)$ points. A coarse-graining of $\lambda$ was performed as a way to estimate the true pdf in order to apply the conformal inference algorithm.
@@
Figure 5: Cdfs for the SIR model obtained from modelling the empirical cdf at different $(\alpha,\beta)$ points with the associated $68\%$ conformal confidence band.
@@
Figure 6: Pdfs for the SIR model obtained from modelling the empirical cdf at different $(\alpha,\beta)$ points with the associated $68\%$ conformal confidence band. The spikes around $\lambda\approx 0.15$ arise from simulated epidemics that die off within a day or so. In a considerably larger sample of simulated epidemics the spikes would not be visible.

### v2

@@ 3.2 Directly Modeling the Empirical CDF
Figure 3: ON/OFF problem: cdfs from modeling the empirical cdf with the latter as the targets at different $(\mu,\nu)$ points, with the associated $68\%$ confidence band computed using conformal inference.
@@
Figure 4: ON/OFF problem: pdfs obtained by differentiating the models of the empirical cdf at different $(\mu,\nu)$ points, with the associated $68\%$ confidence band computed using conformal inference.
@@
Trying to model an intrinsically discrete distribution with a smooth approximation may be beyond an algorithm such as ALFFI and one might argue it is in any case a hopeless task. Nevertheless it is useful to check if an alternative, more direct, approach works better: directly modeling the empirical cdf. At each $\{\mu_{i},\nu_{i}\}$ point, we generate $K=100$ experiments: $\{N_{i}^{(j)}\}_{j=1}^{K}\sim\text{Poiss}(\mu_{i}+\nu_{i})$ and $\{M_{i}^{(j)}\}_{j=1}^{K}\sim\text{Poiss}(\nu_{i})$, for which we generate $K$ instances of the test statistic $\{\lambda_{i}^{(j)}\}_{j=1}^{K}$ according to Eq. 11. These data are used to calculate the empirical cdf $F_{i}(\lambda_{i})$ at every parameter point. The data were then flattened to result in a training set of size $B=10^{7}$, $\mathcal{T}=\{(\mu_{i},\nu_{i},\lambda_{i},F_{i})\}_{i=1}^{B}$.
@@
The cdf was again approximated with a NN with 3 input features $\mathbf{x}=\{\mu,\nu,\lambda\}$ but this time with the targets set to values of the empirical cdf $F$. The architecture consisted of 6 hidden layers with 12 nodes each and a single output. The activation function at each hidden node is a SiLU. The network was trained to minimize the MSE loss with the NAdam optimizer with a fixed learning rate of $3\times 10^{-5}$. The training set is composed of $10^{7}$ examples, which were used in batches of size $60$ for $10^{6}$ iterations, that is, for 6 epochs.
@@
A much better approximation of the cdf is obtained, as shown in Fig. 3 and more importantly, the associated pdfs in Fig. 4 show that the jagged fluctuations in the pdf disappear and we obtain a smooth approximation of the discrete distribution. Modeling the empirical cdf presents yet another advantage: the possibility of using the empirical cdf as the true response variable in the conformal inference algorithm (see Sec. 2.4) and thereby providing a quantification of the uncertainty in the modeling of the cdf and pdf. We use conformal inference to calculate a $68\%$ confidence interval in the cdf space for each value of the test statistic. For the associated pdf, a coarse-graining of $\lambda$ is performed by histogramming $\lambda$ and the coarse-grained $f(\lambda)$ was used as the true response variable (that is, the true pdf). The excellent agreement between the predicted pdf and the histogrammed pdf is shown in Fig. 4 along with the $68\%$ conformal confidence interval at each value of $\lambda$.
3.3 Multistage Modeling of the CDF
@@
Quantile residuals ($\Delta C$): Let $p=\{F_{1},...,F_{k}\}$ be a set of evenly spaced probabilities and let $\hat{F}=\{\hat{F}_{1},...,\hat{F}_{n}\}$ be a set of estimated cdf values at a given $(\mu,\nu)$ parameter point with $n\gg k$. Define $q$ as the quantiles of $\hat{F}$ associated with the probabilities $p$. If we can model the correction function $p=C(q)$ that maps the estimated cdfs $q$ to the exact cdf $p$ with sufficient accuracy then, in principle, the correction function can be used to improve the estimated cdf. Writing $p=q+\Delta C(q)$, we define the quantile residual function as $\Delta C=q-p$ and try to model it as a function of $q$ following the multistage method.
@@
Nevertheless, we wish to see whether a neural network can produce a smooth interpolation of this noisy function by implementing the suggestions described in [28], which are:
  1. 1.
    Use a sine nonlinearity in the first layer and tanh nonlinearity in subsequent layers.
  2. 2.
    Use Glorot weight initialization [35] with a constant factor $\kappa=60$ multiplying the weights in the first layer and no such factor in subsequent layers. It was observed that the $\kappa$ weight factor helps to stabilize the training and the sine nonlinearity helps capture the high frequency.
  3. 3.
    Normalize the target function by the RMS of the outputs of the target function $\Delta C\rightarrow\Delta C/RMS(\Delta C)$, as in Eq. 7.
@@
Further experiments were performed such as adding or omitting dropout, batch normalization, weight regularization, and experimenting with different nonlinear functions such as ReLU, SiLU, tanh, sine, and groupsort [36]. None of these experiments were able to capture the high frequency behavior of the residual target function when trained on the residuals of all parameters in the training data simultaneously. However, we were able to use a simple neural network to model the residual function when modeled point-by-point in the parameter space.
@@
The frequency principle [37] states that neural networks first fit low frequencies and later fit high frequencies during the training process. However, prior studies [38, 39] have demonstrated that in practice standard NNs struggle to fit high frequency functions. We confirm these results and observe that the NNs we used struggle to fit the residual functions even when trained for very long training runs. LSTMs, [40] on the other hand, are known to outperform standard NNs in fitting high frequency and sequential data, where the order of the data is significant [41]. They have been successfully applied in many areas in sequence learning and time series forecasting such as financial forecasting [42], handwriting recognition [43] and speech recognition [44]. In our study, LSTMs achieve reasonably good predictions of the residuals when trained over all $(\mu,\nu)$ values. Although our LSTM models outperform our multi-layer perceptron (MLP) models in this task, the results are unsuitable for use in the MSNN approach due to inadequate precision in the resulting predictions.
@@
Cdf residuals ($\Delta r$): An alternative to modeling the quantile residuals is to model the cdf residuals as a function of $\lambda$, which is closer to what is done in the MSNN approach. At a particular $(\mu,\nu)$ point, the cdf residual is simply the difference between the predicted and the empirical cdf: $\Delta r=\hat{F}-F$. Figure 15 shows that the cdf residuals as a function of $\lambda$ are a slightly smoother function of the parameter points than is $\Delta C$. Nevertheless, the MLP and LSTM models still struggle to capture the high-frequency nature of this residual function. A coarse-graining of $\lambda$ was also performed to reduced the high-frequency of the residual function but to no avail.
@@ 4 Example 2: SIR Model
The lessons learned in the ON/OFF example are applied to modeling the cdf in a different context: epidemiology. We use the SIR model [10, 11] which treats an epidemic as a system in which transitions occur between three states or compartments: susceptible (S), infected (I), and recovered (R). This model has both a stochastic realization as well as a description in terms of ordinary differential equations
@@
where $\theta=\{\alpha,\beta\}$, $x_{n}$ is the observed number of infected individuals at the observation time $t_{n}$ and $I_{n}=I(t_{n},\theta)$ is the predicted mean infected count obtained by solving the system of ordinary differential equations, Eq. 15. We have chosen this example because it demonstrates the practical utility of the methods discussed in this paper for performing inference with intractable statistical models. While all of our major findings from the ON/OFF example apply here, our focus in this section is to improve the modeling of the cdf and to employ additional techniques for uncertainty quantification.
4.1 Directly Modeling the Empirical CDF and Sensitivity Analysis
@@
One of the lessons from the ON/OFF problem is that for discrete distributions the ALFFI algorithm does not yield a sufficiently accurate smooth model of the cdf, which was confirmed in the SIR example. We therefore directly model the empirical cdf as a function of the model parameters $(\alpha,\beta)$ and the test statistic. The training dataset comprises $250$ uniformly sampled $(\alpha,\beta)$ parameter points, where at each $(\alpha_{i},\beta_{i})$ point, $K=400$ epidemics were simulated yielding a set of test statistics $\{\lambda_{i}^{(j)}\}_{j=1}^{K}$ and a set of empirical cdfs $\{F_{i}^{(j)}\}_{j=1}^{K}$. The data were then flattened to result in a training set of size $B\times K=10^{5}$, $\mathcal{T}=\{\alpha_{(i)},\beta_{(i)},\lambda_{(i)},F_{(i)}(\lambda_{(i)})\}_{i=1}^{B\times K}$. All of the models for this example approximate the cdf using a fully-connected neural network with 3 input features $\mathbf{x}=\{\alpha,\beta,\lambda\}$ and the targets the values of the empirical cdfs $F$.
To optimize predictive accuracy, we use the Optuna [45] framework to fine-tune both the neural network architecture and its hyperparameters. Optuna uses a Tree-Structured Parzen Estimator (TPE) [46] algorithm, which is a Bayesian optimization method specifically designed for hyperparameter optimization in machine learning models. The hyperparameter search space is specified in Table 1. 40 optimization trials were executed, where each trial involved sampling a unique neural network configuration and training regimen from the defined search space. For every sampled set of hyperparameters, a neural network was instantiated and trained for 5 epochs. Upon completion of each trial, the model’s performance was evaluated on a validation set, and the best MSE loss achieved was recorded as the objective value.
Algorithm 1 SIR example: data generation, hyperparameter optimization, training and inference
Input: number of parameter points $B$; number of simulated epidemics per point $K$; observed number of infected individuals $D=\{x_{1},...,x_{n}\}$; number of hyperparameter optimization trials $N$; number of training iterations $M$; hyperparameter search space $\Phi$; miscoverage rate $\alpha\in[0,1]$
      Output: learned pdf $\hat{f}(\lambda\mid\alpha,\beta)$ with $1-\alpha$ conformal inference band
1:// data generation
2:Set $\mathcal{T^{\prime}}\leftarrow\emptyset$
3:for $i$ in $\{1,...,B\}$ do
4:     Draw $\alpha_{i}\sim\pi_{\alpha}$, $\beta_{i}\sim\pi_{\beta}$
5:     for $j$ in $\{1,...,K\}$ do
6:         $I_{i}^{(j)}$ = SimulateEpidemic($\alpha_{i},\beta_{i}$) as in Eq. 15
7:         Compute test statistic $\lambda_{i}^{(j)}=\lambda(D;\alpha_{i},\beta_{i},I_{i}^{(j)})$ as in Eq. 16
8:     end for
9:     Compute empirical cdf $F_{i}(\lambda)=F(\lambda|\lambda_{i}^{(1)},...,\lambda_{i}^{(K)})$
10:end for
11:Flatten $\mathcal{T}\leftarrow\mathcal{T}=\{\alpha_{(i)},\beta_{(i)},\lambda_{(i)},F_{(i)}(\lambda_{(i)})\}_{i=1}^{B\times K}$
12:// hyperparameter optimization
13:for $j$ in $\{1,...,N\}$ do
14:     Initialize a candidate vector of hyperparameters $\vec{\phi}_{j}\in\Phi$ and NN model $\mathcal{M}_{\vec{\phi}_{j}}$
15:     Train $\mathcal{M}_{\vec{\phi}_{j}}$ to regress $F$ on $\{\alpha,\beta,\lambda\}$ for 5 epochs
16:     Compute final loss $c_{j}=\mathcal{L}(\mathcal{M}_{\vec{\phi}_{j}},\mathcal{D}_{\text{train}},\mathcal{D}_{\text{valid}})$
17:end for
18:Find the optimal vector of hyperparameters $\vec{\phi}^{*}$ that corresponds to the lowest validation loss $\{c_{j}\}_{j=1}^{N}$
19:// training
20:Initialize model with optimal vector of hyperparameters $\mathcal{M}_{\vec{\phi^{*}}}$
21:Use $\mathcal{T}$ to learn cdf $\hat{F}(\lambda\mid\alpha,\beta)$ via regression of $F$ on $\{\alpha,\beta,\lambda\}$ for $M$ iterations using model $\mathcal{M}_{\vec{\phi^{*}}}$
22:// inference
23:Compute approximated pdf $\hat{f}(\lambda\mid\alpha,\beta)=\frac{\partial\hat{F}(\lambda\mid\alpha,\beta)}{\partial\lambda}$ with autograd
24:Compute $1-\alpha$ conformal inference confidence band as in Eq. 6
25:return learned pdf $\hat{f}(\lambda\mid\alpha,\beta)$ with $1-\alpha$ conformal inference band
@@
The MSE loss is known to be sensitive to outliers in the data. Therefore, we also used the more robust Huber loss [47],
@@
$L_{\delta}(\mathbf{w})=\left\{\begin{array}[]{ll}\frac{1}{2}(y-f(\mathbf{x},\mathbf{w}))^{2}&\text{for }|y-f(\mathbf{x},\mathbf{w})|\leq\delta\\ \delta\left(|y-f(\mathbf{x},\mathbf{w})|-\frac{1}{2}\delta\right),&\text{otherwise,}\end{array}\right.$ (17)
where $\delta$ is a tunable parameter, chosen to be $0.7$. The Huber loss, which is differentiable everywhere and robust to outliers, is an MSE loss for relatively small errors and an absolute loss for larger errors. We compare the performances of the two losses by repeating the hyperparameter optimization procedure outlined above using the Huber loss. The resulting best hyperparameter values for both losses are reported in Table 1 and the hyperparameter importances are shown in Fig. 6. The resulting optimization history for both losses is reported in Fig. 5, which shows that the optimization is robust with respect to the choice of hyperparameters and network architecture.
Our analysis reveals that the mini-batch size is the most influential hyperparameter affecting both loss functions, with larger batch sizes consistently leading to enhanced performance. This observation indicates that, for this particular problem, reducing the stochastic noise in gradient estimations—achieved by increasing the mini-batch size—yields better optimization results. Consequently, using larger mini-batches appears to be beneficial, as it provides more accurate gradient calculations and promotes more stable convergence during training.
Parameter Search space MSE Huber
Number of layers $\in[1,6]$ $3$ $2$
Number of neurons $\in[1,64]$ $55$ $60$
Optimizer [Adam, NAdam, RMSprop, SGD] RMSprop Adam
Learning rate $\in[10^{-6},10^{-2}]$ $0.00268$ $0.00543$
Batch size $\in[50,3\times 10^{4}]$ $19,980$ $19,961$
Activation [ReLU, LeakyReLU, SELU, PReLU] LeakyReLU PReLU
Table 1: Hyperparameter tuning ranges and best values. The right two columns show the best values from the optimization process for the MSE and Huber losses, respectively.
Figure 5: Optimization history using MSE loss (top) and using Huber loss (bottom). For each trial, the blue dots represent the objective value. The red line shows the best objective value obtained up to a given trial.
Figure 6: Hyperparameter importances using MSE loss (top) and using Huber loss (bottom).
Following the optimization process for our model hyperparameters, we initialized an MLP with the optimal hyperparameters identified for the MSE loss function. This network was then subjected to an extended training regimen of $2\times 10^{5}$ iterations, equivalent to 39,960 epochs. Throughout training, we saved the model that achieved the lowest validation loss. Figures 8 and 9 depict the resulting estimated cdfs and pdfs with the 68% conformal confidence intervals at different values of $\lambda$ and for different $(\alpha,\beta)$ points. To estimate the true pdf, which is necessary for the conformal inference algorithm, we performed a coarse-graining of $\lambda$. This entire procedure was repeated using a model trained with the Huber loss, using the optimal hyperparameters determined for the Huber loss. Figure 7 presents the training and validation loss curves over the course of training for both models, indicating that the Huber loss model achieves lower training and validation losses compared to the MSE loss model. Figures 10 and 11 display the resulting estimated cdfs and pdfs with the 68% conformal confidence intervals at different values of $\lambda$ and for different $(\alpha,\beta)$ points. Pseudocode for the data generation, training, hyperparameter optimization, training and inference is shown in Algorithm 1.
Figure 7: SIR problem: training (left) and validation (right) loss curves.
@@
Figure 8: Cdfs for the optimized SIR model (MSE loss) obtained from modeling the empirical cdf at different $(\alpha,\beta)$ points with the associated $68\%$ conformal confidence band.
@@
Figure 9: Pdfs for the optimized SIR model (MSE loss) obtained from modeling the empirical cdf at different $(\alpha,\beta)$ points with the associated $68\%$ conformal confidence band. The spikes around $\lambda\approx 0.15$ arise from simulated epidemics that die off within a day or so of the start of the simulated epidemics. In a considerably larger sample of simulated epidemics the spikes would not be visible.
Figure 10: Cdfs for the optimized SIR model (Huber loss) obtained from modeling the empirical cdf at different $(\alpha,\beta)$ points with the associated $68\%$ conformal confidence band.
Figure 11: Pdfs for the optimized SIR model (Huber loss) obtained from modeling the empirical cdf at different $(\alpha,\beta)$ points with the associated $68\%$ conformal confidence band. The spikes around $\lambda\approx 0.15$ arise from simulated epidemics that die off within a day or so of the start of the simulated epidemics. In a considerably larger sample of simulated epidemics the spikes would not be visible.